every year, the assigned teachers determine the body mass index of students. in a certain public junior high school, a study finds that 10% of grade 7 students observe are underweight. A sample of 780 Grade 7 students was randomly chosen and it was found out that 125 of them are underweight. Is this claim different for their grade level age? Use 0.05 level of significance


1
Expert's answer
2021-06-27T17:24:58-0400
"\hat{p}=\dfrac{125}{780}=\dfrac{25}{156}"

The following null and alternative hypotheses for the population proportion needs to be tested:

"H_0:p=0.1"

"H_1:p\not=0.1"

This corresponds to a two-tailed test, for which a z-test for one population proportion will be used.

Based on the information provided, the significance level is "\alpha=0.05," and the critical value for a two-tailed test is "z_c=1.96."

The rejection region for this two-tailed test is "R=\{z:|z|>1.96\}".

The z-statistic is computed as follows:


"z=\dfrac{\hat{p}-p_0}{\sqrt{\dfrac{p_0(1-p_0)}{n}}}"

"=\dfrac{\dfrac{25}{156}-0.1}{\sqrt{\dfrac{0.1(1-0.1)}{780}}}\approx5.61"

Since it is observed that "|z|=5.61>1.96=z_c," it is then concluded that the null hypothesis is rejected. Therefore, there is enough evidence to claim that the population proportion "p" is different than "0.1," at the "\alpha=0.05" significance level.


Using the P-value approach: The p-value is "p=2P(z>5.61)=0," and since "p=0<0.05=\alpha,"

it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population proportion "p" is different than "0.1," at the "\alpha=0.05" significance level.